OpenStudy (anonymous):

What is the slope of a line perpendicular to x - 5y = -10?

OpenStudy (diyadiya):

Write it in the form y=mx+b where m is the slope :)

OpenStudy (anonymous):

Im confused so would u write it 5y=mx-10?

OpenStudy (diyadiya):

Nope

OpenStudy (diyadiya):

First subtract x from both sides

OpenStudy (anonymous):

so -5y=-10-x?

OpenStudy (diyadiya):

Right Now we need this in the form y=mx+c

OpenStudy (diyadiya):

just rearrange it and you'll get -5y=-x-10 okay?

OpenStudy (anonymous):

okay

OpenStudy (diyadiya):

Now you don't need that -5 on the left side ,do you?

OpenStudy (anonymous):

no

OpenStudy (diyadiya):

Ok so what do you do?

OpenStudy (anonymous):

divide ?

OpenStudy (diyadiya):

Yes! divide both sides by -5

OpenStudy (diyadiya):

and be carefull when you divide the right side ,you should divide each term

OpenStudy (anonymous):

y=5x+2 ?

OpenStudy (diyadiya):

Nope

OpenStudy (diyadiya):

y=(-x/-5)- (10/-5)

OpenStudy (anonymous):

so m= 2

OpenStudy (diyadiya):

Nope

OpenStudy (diyadiya):

\[y= \frac{-x}{-5}- \frac{10}{-5}\]

OpenStudy (anonymous):

-5x+2

OpenStudy (diyadiya):

\[y= \frac{x}{5}-2\]

OpenStudy (diyadiya):

\[y= \frac{1}{5}x-2\]

OpenStudy (diyadiya):

this is of the form y=mx+c So m is?

OpenStudy (anonymous):

-2

OpenStudy (diyadiya):

\[y= \frac{1}{5}x-2\]\[y=m \ x +c\] Now tell me what is m ?

OpenStudy (diyadiya):

just compare both equation

OpenStudy (anonymous):

1/5 but thats not one of my anwsers...

OpenStudy (diyadiya):

0.2?

OpenStudy (anonymous):

the anwsers are 2 -3 -4 -5 -2

OpenStudy (diyadiya):

Sorry for the wrong answer =|

OpenStudy (anonymous):

Im confused. idk what to put. i see that we did it right, but i dont understand how the answer doesnt come out to be one of those...

OpenStudy (apoorvk):

okay.. so for a line of the standard slope intercept form "y=mx+c" slope is 'm' and for a line perpendicular to this one, its slope is automatically "-1/m". so can you rearrange the above line's equation above in the form above, and find 'm'?? post what you get.

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

this just confuses me....

OpenStudy (diyadiya):

OHHH YEH its PERPENDICULAR!!!! OOPS

OpenStudy (diyadiya):

Its just the same when its perpendicular the equation becomes \[y=\frac{-1}{m}x+c\]

OpenStudy (anonymous):

m=-5 and c= -10?

OpenStudy (apoorvk):

let me show you S T E P - B Y - S T E P. V E R Y S L O W L Y! the line is x-5y = -10 or, 5y= x + 10 y = x/5 + 2 now comparing (i) with y=mx+c. coefficient of x = m = slope = 1/5 so slope of the perpendicular line = -1/m = -5 Clear?

OpenStudy (anonymous):

some what

OpenStudy (apoorvk):

go take a break, and come back to this problem with a clear mind. it 'll come to you naturally then :) don't fret